# CLUENet: Cluster Attention Makes Neural Networks Have Eyes

### v1

@@ Abstract
Despite the success of convolution- and attention-based models in vision tasks, their rigid receptive fields and complex architectures limit their ability to model irregular spatial patterns and hinder interpretability, therefore posing challenges for tasks requiring high model transparency. Clustering paradigms offer promising interpretability and flexible semantic modeling, but suffer from limited accuracy, low efficiency, and gradient vanishing during training. To address these issues, we propose CLUster attEntion Network (CLUENet), an transparent deep architecture for visual semantic understanding. We propose three key innovations include (i) a Global Soft Aggregation and Hard Assignment with a Temperature-Scaled Cosin Attention and gated residual connections for enhanced local modeling, (ii) inter-block Hard and Shared Feature Dispatching, and (iii) an improved cluster pooling strategy. These enhancements significantly improve both classification performance and visual interpretability. Experiments on CIFAR-100 and Mini-ImageNet demonstrate that CLUENet outperforms existing clustering methods and mainstream visual models, offering a compelling balance of accuracy, efficiency, and transparency.
@@ Introduction
Deep Neural Networks have become the primary approach for image analysis, and various deep architectures have been proposed and differ in their modeling of intrinsic image structures Early deep learning methods predominantly relied on the convolutional paradigm (ref1; ref2; ref3; ref4; ref5; ref6; ref7). These Convolutional Neural Networks (CNNs) extracts features by applying shared-weight kernels locally via sliding windows, inherently offering translation equivariance and local context modeling (ref1). With the rise of Transformer, the attention-based paradigm (ref8; ref9; ref10; ref11; ref12) has emerged as a powerful alternative for image analysis. Transformers leverage global modeling capabilities to extract features through adaptive weight allocation, however, this comes at the cost of significant computational complexity.
@@
Despite significant progress in vision tasks, deep networks remain limited in modeling the workings of human visual system (ref13; ref14; ref15). Unlike artificial neural network systems, the human visual system dynamically clusters scenes by grouping semantically consistent parts (ref16; ref17; ref18). Conversely, convolutional and attention paradigms rely on fixed geometric windows (convolutional kernels or patches), constraining their ability to adaptively capture irregular object in images (ref19). Moreover, both paradigms prioritize performance through escalating architectural complexity, generating opaque abstractions that lack human interpretability  (ref20).
@@
To bridge this gap, the clustering paradigm (ref16; ref19; ref21) has emerged as a transparent alternative inspired by clustering algorithms. This approach explicitly models images as sets of semantically coherent pixel groups, directly mirroring human perceptual organization. Pioneering this direction, Ma et al. (ref21) introduced Context-Cluster (CoC) for general vision tasks, leveraging unsupervised clustering at multiple feature levels to enable intrinsic interpretability via cluster visualization. Building on this, Chen et al. (ref19) proposed Feature Extraction with Clustering (FEC) which integrated clustering into pooling layers, enhancing model performance and enabling receptive field tracing for each feature point. ClusterFormer (ref16) incorporated cross-attention within iterative Expectation-Maximization (EM) steps to refine initial clusters based on adjacency, enhancing cluster initialization and expanding spatial receptive fields.
@@
Despite the enhanced model transparency offered by clustering paradigms, three key challenges persist: (1) Suboptimal performance: Current clustering-based models underperform relative to state-of-the-art vision architectures (ref5; ref11), typically exceeding only pre-2022 baselines. (2) Limited receptive field: Computational constraints force methods like CoC and ClusterFormer to perform local clustering within isolated windows. This design lacks inter-window communication, compromising semantic continuity and feature quality. (3) Gradient vanishing in specialized components: FEC’s similarity projection layer in cluster pooling suffers gradient vanishing on small datasets, freezing during training and causing redundancy and inefficiency.
@@
In this paper, we propose a novel CLUster attEtion Network (CLUENet) to address the above discussed limitations. The contributions of CLUENet is summarized as follows:
@@
  • Global Soft Aggregation and Hard Assignment. CLUENet computes global similarities between cluster centers and all pixels to form soft clusters via weighted fusion. A gated residual module supplements local context, and a global query head enables precise hard assignment for each pixel.
@@
  • Efficient Aggregation with Shared Assignment. We employ cosine attention with learnable temperature in half-precision via FlashAttention for fast, memory-efficient aggregation. Hard assignment matrices are shared across blocks within each stage to reduce redundancy and enhance stability.
@@
  • Enhanced Cluster Pooling. We perform clustering and pooling in similarity space and project results back to feature space via a perceptron, effectively alleviating gradient vanishing and enhancing performance.
@@
Model interpretability. Deep Neural Networks (DNNs) have achieved remarkable success in computer vision, yet their black-box nature raises concerns, especially in safety-critical domains such as medical diagnosis and autonomous driving (ref27). Convolutional Neural Networks (CNNs) inherently lack intrinsic interpretability, making it difficult to intuitively reveal their decision rationale. Model-inspired deep architectures (huang2020learning; huang2022winnet; huang2025lightweight; pu2022mixed) improve model interpretability by embedding the formation model and priors into the model. Existing visualization methods mainly focus on class activation maps (ref28), feature visualization (ref29), and input sensitivity analysis (ref30). While these approaches improve interpretability to some extent, they are largely post-hoc and do not fully reveal the true internal decision mechanisms. With the rise of Vision Transformers (ViTs), attention-based interpretability has become a new research hotspot, including attention visualization (ref8) and emergent segmentation phenomena in self-supervised learning (ref26). However, the interpretability of ViTs remains limited and uncertain. On the one hand, attention weights may be influenced by input noise or randomness during training, questioning their reliability as explanations (ref31). On the other hand, most ViT interpretability still relies on post-hoc methods rather than uncovering internal decision processes. Recently, research focus has been shifting towards enhancing model transparency through intrinsically interpretable model architectures. Sparse Transformers (ref32) introduces sparse attention to highlight decision-critical features. Deep Nearest Centroids (ref33) offers analogy-based explanations by learning feature-to-centroid distances. And white-box Transformers like CRATE (yu2023white) designed for observable and traceable attention paths and feature flows.
@@
Clustering-based Model. The clustering paradigm proposed in 2023 (ref21) innovatively applies the Simple Linear Iterative Clustering (ref35) to general visual representation tasks. This approach achieves strong performance in classification, detection, segmentation, and 3D point cloud processing while generating semantic grouping and visual explanations during inference, offering intuitive decision cues for users. As the first systematic clustering paradigm model, CoC (ref21) represents images as spatially structured pixel groups and extract features through a deep architecture inspired by a simplified clustering algorithm. Recently, ClusterFormer (ref16) integrates clustering with cross-attention in iterative EM steps. However, it relies on local windows, limiting its receptive field. FEC (ref19) supports global clustering but simplifies feature aggregation by averaging pixels per cluster without similarity weighting, reducing memory usage but risking gradient vanishing and limiting semantic complexity due to fewer clusters. Other recent works extend clustering and semantic aggregation to broader settings, highlighting the generality of cluster-based semantic modeling (ju2023glcc; liang2025from; guo2024depth).
@@
Figure 2: (a) Overall architecture of the CLUENet with a four-stage pyramid network. (b) The key components within the Global Feature Clustering (GFC) block, illustrating Global and Soft Feature Aggregation (GSFA) that updates cluster centers from all pixels, and Hard and Shared Feature Dispatching (HSFD) that updates pixel features according to their assigned cluster centers; (c) The Improved Cluster Pooling (ICP) block, depicting how pixel features are grouped into clusters in similarity space while preserving hierarchical structure
@@
In this paper, we present CLUster attEntion Network (CLUENet), an transparent deep architecture for visual semantic understanding built on the clustering paradigm. As shown the overview of CLUENet in Fig. 2 (a), the proposed CLUENet uses a four-stage pyramid structure and consists of three novel components: (i) A Positional-aware Feature Embedding (PFE) Block encodes spatial relationships while preserving semantic locality. (ii) Global Feature Clustering (GFC) Blocks (see Fig. 2(b)) enable interpretable and flexible semantic modeling through global-soft aggregation paired with efficient hard dispatching overcoming the fixed-window limiations. (iii) Improved Cluster Pooling (ICP) Blocks (see Fig. 2(c)) perform semantic information aggregation by grouping pixels into clusters in similarity space, maintaining hierarchical structure while while mitigating gradient vanishing. In the following, we introduce the key components of CLUENet in details.
@@
For visual semantic clustering, positional information plays a crucial role in distinguishing structured visual regions. Conventional methods rely on fixed grid coordinates for positional embedding, however, this rigid approach lack adaptability to scale variations. This limitation becomes particularly detrimental in scale-sensitive tasks such as instance and semantic segmentation (ref36).
@@
Figure 3: The detials of the Global Feature Clustering (GFC) block. Global and Soft Feature Aggregation (GSFA) inclues center initialization, Temperature-Scaled Cosine Attention, and Gated Fusion Mechanism. Hard and Shared Feature Dispatching (HSFD) includes multi-head query projection, hard clustering, and assignment shared across blocks.
@@
The GFC block serves as the core visual semantic clustering component in CLUENet. As shown in Fig. 3, the GFC block integrates a Global and Soft Feature Aggregation (FSFA) for cluster centers updating and a Hard and Shared Feature Dispatching (HSFD) for feature updating. They jointy enable effectively semantic relationships capturing while maintaining high computational efficiency.
@@
Global aggregation enables each cluster center to dynamically attend to all pixels across the image, capturing long-range dependencies beyond local neighborhoods and promoting continuous clusters instead of disjoint window-based fragmented segments. And soft aggregation provides flexible, weighted attention over multiple pixels, enhancing feature grouping and improving training stability.
@@
The cluster centers and pixel features are reshaped into 1D sequences $\mathbf{c}_{i}\in\mathbb{R}^{m\times d^{\prime}},\mathbf{p}_{i}\in\mathbb{R}^{n\times d^{\prime}},\text{with }n=H\times W.$
@@
where ${{\mathbf{S}}}_{C}\in\mathbb{R}^{m\times n}$ is the similarity matrix between cluster centers and all pixels, $\mathbf{c}^{\prime}_{v}\in\mathbb{R}^{m\times d^{\prime}}$ represents the soft cluster centers obtained by weighted aggregation, $\tau$ is a learnable temperature parameter.
This design allows each cluster center to retain clustering weights for all pixels across the global field while emphasizing differences in similarity, enabling selective and nuanced aggregation.
@@
where $\sigma(\cdot)$ is the Sigmoid function, $f(\cdot)$ represents the gating network implemented as a two-layer perceptron, $\mathbf{g}\in[0,1]^{m}$ is the gating weight vector which is broadcasted along the feature dimension.
@@
A hard assignment strategy is then applied, that is each pixel selects only its most similar cluster center. The resulting sparse similarity matrix $\widetilde{\mathbf{S}}_{P}$ is reused in the subsequent clustering attention module:
@@
where $\mathbf{S}_{P}\in\mathbb{R}^{n\times m}$ denotes the similarity between each pixel and the query heads, and $\widetilde{\mathbf{S}}_{P}$ is the sparsified matrix after hard assignment, FC($\cdot$) denotes a fully connected layer used to project features from hidden dimension $d^{\prime}$ back to the original dimension $d$.
Multi-Head Clustering: Inspired by multi-head attention (ref23), we split semantic queries and pixel features along the feature dimension into $M$ heads, enabling parallel cluster assignment and improving semantic modeling accuracy. Specifically, $\mathbf{p}_{i}$ is first partitioned into $M$ heads $\mathbf{p}_{i}=\{\mathbf{p}_{s}^{(m)}\in\mathbb{R}^{n\times d^{\prime}/M}\}_{i=1}^{M},i\in\{s,v\}$. These multi-head pixel features induce multi-head cluster centers in the value space $\widetilde{\mathbf{C}}_{v}=\{\widetilde{\mathbf{C}}_{v}^{(m)}\in\mathbb{R}^{m\times d^{\prime}/M}\}_{i=1}^{M}$. Then, the multi-head cluster centers $\widetilde{\mathbf{C}}_{v}$ are concatenated along the channel dimension and linearly projected via a learnable matrix $\mathbf{W}_{q}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ to form query features $\mathbf{Q}^{\prime}=\mathbf{W}_{q}\cdot\mathrm{concat}\left(\widetilde{\mathbf{C}}_{v}^{(1)},\ldots,\widetilde{\mathbf{C}}_{v}^{(M)}\right)\in\mathbb{R}^{m\times d^{\prime}}$. The projected queries $\mathbf{q}^{\prime}$ are also split back into $M$ heads $\mathbf{q}=\{\mathbf{q}^{(m)}\in\mathbb{R}^{m\times d^{\prime}/M}\}_{i=1}^{M}$ which serve as semantic query heads for the hard assignment in each subspace.
@@
We identify a key limitation in the cluster pooling design of FEC (ref19). As shown in Fig. 4, the projection layer $\operatorname{proj_{f}}$ which maps features into the similarity space, merely marks pixel-cluster associations during forward pass but does not participate in the actual aggregation. This leads to a disconnection between the forward data flow and the learnable parameters of $\operatorname{proj_{f}}$, resulting in ineffective gradient backpropagation and a near-identity behavior during training.
To address this, we propose to perform average pooling directly in the similarity space, making the aggregation explicitly dependent on $\operatorname{proj_{f}}$’s output and ensuring it is learning. Meanwhile, the $\operatorname{proj_{v}}$ mapping layer is redefined to map the aggregated similarity information back to the feature space. Notably, we adopt a two-layer perceptron for $\operatorname{proj_{v}}$ instead of a linear layer, as experiments show that single-layer designs offer limited performance gain. With these modifications, both $\operatorname{proj_{f}}$ and $\operatorname{proj_{v}}$ are effectively optimized during training, leading to significant performance improvements.
Figure 4: Cluster pooling configurations. (a) In FEC, $\operatorname{proj_{f}}$ only guides pixel selection and does not update, (b) connecting $\operatorname{proj_{f}}$ and $\operatorname{proj_{v}}$ avoids gradient issues but get no performance, (c) The proposed cluster pooling performs in similarity space, enabling effective training and improved performance.
@@
This section presents a comprehensive evaluation of the proposed CLUENet, conducts visualiation analysis and ablation studies [^1] Further experimental details, visualizations, and failure case analyses are available in the supplementary material..
@@
ShuffleNet_v2 (x1.5) (ref6) 2.58 0.31 78.39 90.40 1194.72±23.26 1.1
ShuffleNet_v2 (x2.0) (ref6) 6.52 0.64 79.63 90.93 1205.17±7.59 1.3
ConvNetxt_v2 (A) (ref43) 3.42 0.55 71.15 84.97 1172.04±34.80 1.6
ConvNetxt_v2 (F) (ref43) 4.89 0.79 73.14 86.33 1177.66±21.15 1.8
ConvNetxt_v2 (N) (ref43) 15.05 2.46 75.04 87.48 1148.19±28.45 2.8
Attention-based method PVT_v2 (b0) (ref11) 3.44 0.54 75.34 88.52 1195.32±31.63 1.7
PVT_v2 (b1) (ref11) 13.55 2.06 77.57 89.86 1154.58±39.86 2.9
@@
Datasets. CIFAR-100 and Mini-ImageNet are used for experiment. CIFAR-100 contains 50K training and 10K validation images of size $32\times 32$. Mini-ImageNet is a subset of ImageNet-1K with 100 classes and 600 images per class. We split Mini-ImageNet 4:1 into training and validation sets, ensuring balanced class distribution.
We follow the standard data augmentation procedure during training. For CIFAR-100, random horizontal flip and normalization is applied. For Mini-ImageNet, random crop, horizontal flip, color jitter, random erase, and normalization are applied for training.
Training Details. All models are trained using the AdamW optimizer with momentum 0.9 and cosine decay. Weight decay is set to 0.05. The learning rate is auto-tuned using findLR (ref42), with a warmup of 5 epochs. The batch size for CIFAR-100 is 256, and the batch size for Mini-ImageNet is 128. All experiments were conducted on a computer with a NVIDIA RTX 4090 (24GB) GPU.
Evaluation Protocol. Following common practice (ref16; ref19; ref21), all models are evaluated on the validation set after completing 100 training epochs, rather than using the model with the best validation performance, to ensure a fair comparison. Each model is evaluated 7 times, the first 2 runs are discarded to avoid cold-start effects. We report the average of the last 5 runs for classification accuracy (Top-1 and Top-3), throughput (Frame Per Second, FPS), number of parameters (#Param), FLOPs, and inference memory usage, while all metrics except FPS show no significant change.

### v2

@@ Abstract
Despite the success of convolution- and attention-based models in vision tasks, their rigid receptive fields and complex architectures limit their ability to model irregular spatial patterns and hinder interpretability, thereby posing challenges for tasks requiring high model transparency. Clustering paradigms offer promising interpretability and flexible semantic modeling, but suffer from limited accuracy, low efficiency, and gradient vanishing during training. To address these issues, we propose the CLUster attEntion Network (CLUENet), a transparent deep architecture for visual semantic understanding. Specifically, we introduce three key innovations, including (i) a Global and Soft Feature Aggregation with a Temperature-Scaled Cosine Attention for capturing long-range dependencies and a Gated Fusion Mechanism for enhanced local modeling, (ii) Hard and Shared Feature Dispatching, and (iii) an Improved Cluster Pooling Block. These enhancements significantly improve both classification performance and visual interpretability. Experiments on CIFAR-100 and Mini-ImageNet demonstrate that CLUENet outperforms existing clustering methods and mainstream visual models, offering a compelling balance of accuracy, efficiency, and transparency.
@@ Introduction
Deep Neural Networks have become the primary approach for image analysis, and various deep architectures have been proposed and differ in their modeling of intrinsic image structures. Early deep learning methods predominantly relied on the convolutional paradigm (ref1; ref2; ref3; ref4; ref5; ref6; ref7). These Convolutional Neural Networks (CNNs) extract features by applying shared-weight kernels locally via sliding windows, inherently offering translation equivariance and local context modeling (ref1). With the rise of Transformers, the attention-based paradigm (ref8; ref9; ref10; ref11; ref12) has emerged as a powerful alternative for image analysis. Transformers leverage global modeling capabilities to extract features through adaptive weight allocation. However, this comes at the cost of significant computational complexity.
@@
Despite significant progress in vision tasks, deep networks remain limited in modeling the workings of human visual system (ref13; ref14; ref15). Unlike artificial neural network systems, the human visual system dynamically clusters scenes by grouping semantically consistent parts (ref16; ref17; ref18). Conversely, convolutional and attention paradigms rely on fixed geometric windows (convolutional kernels or patches), constraining their ability to adaptively capture irregular objects in images (ref19). Moreover, both paradigms prioritize performance through escalating architectural complexity, generating opaque abstractions that lack human interpretability  (ref20).
@@
To bridge this gap, the clustering paradigm (ref16; ref19; ref21) has emerged as a transparent alternative inspired by clustering algorithms. This approach explicitly models images as sets of semantically coherent pixel groups, directly mirroring human perceptual organization. Pioneering this direction, Ma et al. (ref21) introduced Context-Cluster (CoC) for general vision tasks, leveraging unsupervised clustering at multiple feature levels to enable intrinsic interpretability via cluster visualization. Building on this, Chen et al. (ref19) proposed Feature Extraction with Clustering (FEC) which integrated clustering into pooling layers, enhancing model performance and enabling receptive field tracing for each feature point. ClusterFormer (ref16) incorporated cross-attention within iterative Expectation-Maximization (EM) steps to refine initial clusters solely from adjacency, enhancing cluster initialization and expanding spatial receptive fields.
@@
Despite the enhanced model transparency offered by clustering paradigms, three key challenges persist: (1) Suboptimal performance: Current clustering-based models underperform relative to state-of-the-art vision architectures (ref5; ref11), typically exceeding only pre-2022 baselines. (2) Limited receptive field: Computational constraints force methods like CoC and ClusterFormer to perform local clustering within isolated windows. This design lacks inter-window communication, compromising semantic continuity and feature quality. (3) Gradient vanishing in specialized components: FEC’s similarity projection layer in cluster pooling suffers from gradient vanishing on small datasets, freezing during training and causing redundancy and inefficiency.
@@
In this paper, we propose a novel CLUster attEtion Network (CLUENet) to address the above discussed limitations. The contributions of CLUENet are summarized as follows:
@@
  • Global Soft Aggregation and Hard Assignment allows CLUENet to compute global similarities between cluster centers and all pixels to form soft clusters via weighted fusion, while incorporating a gated residual module to supplement local context and a global query head to enable precise hard assignment for each pixel.
@@
  • Efficient Aggregation with Shared Assignment employs cosine attention with learnable temperature in half-precision via FlashAttention for fast, memory-efficient aggregation, and shares hard assignment matrices across blocks within each stage to reduce redundancy and enhance stability.
@@
  • Improved Cluster Pooling performs clustering and pooling in similarity space and projects results back to feature space via a perceptron, effectively alleviating gradient vanishing and enhancing performance.
@@
Model interpretability. Deep Neural Networks (DNNs) have achieved remarkable success in computer vision, yet their black-box nature raises concerns, especially in safety-critical domains such as medical diagnosis and autonomous driving (ref27). Convolutional Neural Networks (CNNs) inherently lack intrinsic interpretability, making it difficult to intuitively reveal their decision rationale. Model-inspired deep architectures (huang2020learning; huang2022winnet; huang2025lightweight; pu2022mixed) improve model interpretability by embedding the formation model and priors into the model. Existing visualization methods mainly focus on class activation maps (ref28), feature visualization (ref29), and input sensitivity analysis (ref30). While these approaches improve interpretability to some extent, they are largely post-hoc and do not fully reveal the true internal decision mechanisms. With the rise of Vision Transformers (ViTs), attention-based interpretability has become a new research hotspot, including attention visualization (ref8) and emergent segmentation phenomena in self-supervised learning (ref26). However, the interpretability of ViTs remains limited and uncertain. On the one hand, attention weights may be influenced by input noise or randomness during training, questioning their reliability as explanations (ref31). On the other hand, most ViT interpretability still relies on post-hoc methods rather than uncovering internal decision processes. Recently, research focus has been shifting towards enhancing model transparency through intrinsically interpretable model architectures. Sparse Transformers (ref32) introduce sparse attention to highlight decision-critical features. Deep Nearest Centroids (ref33) offers analogy-based explanations by learning feature-to-centroid distances. And white-box Transformers like CRATE (yu2023white) are designed for observable and traceable attention paths and feature flows.
@@
Clustering-based Model. The clustering paradigm proposed in 2023 (ref21) innovatively applies the Simple Linear Iterative Clustering (ref35) to general visual representation tasks. This approach achieves strong performance in classification, detection, segmentation, and 3D point cloud processing while generating semantic grouping and visual explanations during inference, offering intuitive decision cues for users. As the first systematic clustering paradigm model, CoC (ref21) represents images as spatially structured pixel groups and extracts features through a deep architecture inspired by a simplified clustering algorithm. Recently, ClusterFormer (ref16) integrates clustering with cross-attention in iterative EM steps. However, it relies on local windows, limiting its receptive field. FEC (ref19) supports global clustering but simplifies feature aggregation by averaging pixels per cluster without similarity weighting, reducing memory usage but risking gradient vanishing and limiting semantic complexity due to fewer clusters. Other recent works extend clustering and semantic aggregation to broader settings, highlighting the generality of cluster-based semantic modeling (ju2023glcc; liang2025from; guo2024depth).
@@
Figure 2: (a) Overall architecture of the CLUENet with a four-stage pyramid network; (b) The key components within the Global Feature Clustering (GFC) block, illustrating Global and Soft Feature Aggregation (GSFA) that updates cluster centers from all pixels, and Hard and Shared Feature Dispatching (HSFD) that updates pixel features according to their assigned cluster centers; (c) The Improved Cluster Pooling (ICP) block, depicting how pixel features are grouped into clusters in similarity space while preserving hierarchical structure.
@@
In this paper, we present CLUster attEntion Network (CLUENet), a transparent deep architecture for visual semantic understanding built on the clustering paradigm. As shown in the overview of CLUENet in Fig. 2(a), the proposed CLUENet uses a four-stage pyramid structure and consists of three novel components: (i) A Positional-aware Feature Embedding (PFE) Block encodes spatial relationships while preserving semantic locality. (ii) Global Feature Clustering (GFC) Blocks (see Fig. 2(b)) enable interpretable and flexible semantic modeling through global-soft aggregation paired with efficient hard dispatching, overcoming the fixed-window limitations. (iii) Improved Cluster Pooling (ICP) Blocks (see Fig. 2(c)) perform semantic information aggregation by grouping pixels into clusters in similarity space, maintaining hierarchical structure while mitigating gradient vanishing. In the following, we introduce the key components of CLUENet in detail.
@@
For visual semantic clustering, positional information plays a crucial role in distinguishing structured visual regions. Conventional methods rely on fixed grid coordinates for positional embedding. However, this rigid approach lacks adaptability to scale variations. This limitation becomes particularly detrimental in scale-sensitive tasks such as instance and semantic segmentation (ref36).
@@
Figure 3: The details of the Global Feature Clustering (GFC) block. Global and Soft Feature Aggregation (GSFA) includes center initialization, Temperature-Scaled Cosine Attention, and Gated Fusion Mechanism. Hard and Shared Feature Dispatching (HSFD) includes multi-head query projection, hard clustering, and assignment shared across blocks.
@@
The GFC block serves as the core visual semantic clustering component in CLUENet. As shown in Fig. 3, the GFC block integrates a Global and Soft Feature Aggregation (GSFA) for cluster center updating and a Hard and Shared Feature Dispatching (HSFD) for pixel feature updating. They jointly enable the effective capture of semantic relationships while maintaining high computational efficiency.
@@
Global aggregation enables each cluster center to dynamically attend to all pixels across the image, capturing long-range dependencies beyond local neighborhoods and promoting continuous clusters instead of disjoint window-based fragmented segments. And soft aggregation provides flexible, weighted attention over multiple pixels, enhancing feature aggregation and improving training stability.
@@
The cluster centers and pixel features are reshaped into 1D sequences $\mathbf{c}_{i}\in\mathbb{R}^{m\times d^{\prime}},\mathbf{p}_{i}\in\mathbb{R}^{n\times d^{\prime}},\text{where }n=H\times W.$
@@
where ${{\mathbf{S}}}_{C}\in\mathbb{R}^{m\times n}$ is the similarity matrix between cluster centers and all pixels, $\mathbf{c}^{\prime}_{v}\in\mathbb{R}^{m\times d^{\prime}}$ represents the soft cluster centers obtained by weighted aggregation, and $\tau$ is a learnable temperature parameter.
This design allows each cluster center to retain similarity weights for all pixels across the global field while emphasizing differences in similarity, enabling selective and nuanced aggregation.
@@
where $\sigma(\cdot)$ is the sigmoid function, $f(\cdot)$ represents the gating network implemented as a two-layer perceptron, and $\mathbf{g}\in[0,1]^{m}$ is the gating weight vector which is broadcast along the feature dimension.
@@
A hard assignment strategy is then applied, where each pixel selects only its most similar cluster center. The resulting sparse similarity matrix $\widetilde{\mathbf{S}}_{P}$ is reused in the subsequent blocks:
@@
where $\mathbf{S}_{P}\in\mathbb{R}^{n\times m}$ denotes the similarity between each pixel and the query heads, $\widetilde{\mathbf{S}}_{P}$ is the sparsified matrix after hard assignment, and FC($\cdot$) denotes a fully connected layer used to project features from hidden dimension $d^{\prime}$ back to the original dimension $d$.
Multi-Head Clustering: Inspired by multi-head attention (ref23), we split semantic queries and pixel features along the feature dimension into $M$ heads, enabling parallel cluster assignment and improving semantic modeling accuracy. Specifically, $\mathbf{p}_{i}$ is first partitioned into $M$ heads $\mathbf{p}_{i}=\{\mathbf{p}_{i}^{(k)}\in\mathbb{R}^{n\times d^{\prime}/M}\}_{k=1}^{M},i\in\{s,v\}$. These multi-head pixel features induce multi-head cluster centers in the value space $\widetilde{\mathbf{C}}_{v}=\{\widetilde{\mathbf{C}}_{v}^{(k)}\in\mathbb{R}^{m\times d^{\prime}/M}\}_{k=1}^{M}$. Then, the multi-head cluster centers $\widetilde{\mathbf{C}}_{v}$ are concatenated along the channel dimension and linearly projected via a learnable matrix $\mathbf{W}_{q}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ to form query features $\mathbf{q}^{\prime}=\mathrm{concat}\left(\widetilde{\mathbf{C}}_{v}^{(1)},\ldots,\widetilde{\mathbf{C}}_{v}^{(M)}\right)\cdot\mathbf{W}_{q}\in\mathbb{R}^{m\times d^{\prime}}$. The projected queries $\mathbf{q}^{\prime}$ are also split back into $M$ heads $\mathbf{q}=\{\mathbf{q}^{(k)}\in\mathbb{R}^{m\times d^{\prime}/M}\}_{k=1}^{M}$ which serve as semantic query heads for the hard assignment in each subspace.
@@
We identify a key limitation in the cluster pooling design of FEC (ref19). As shown in Fig. 4, the projection layer $\operatorname{proj_{f}}$ which maps features into the similarity space, merely marks pixel-cluster associations during the forward pass but does not participate in the actual aggregation. This leads to a disconnection between the forward data flow and the learnable parameters of $\operatorname{proj_{f}}$, resulting in ineffective gradient backpropagation and a near-identity behavior during training.
To address this, we propose to perform average pooling directly in the similarity space, making the aggregation explicitly dependent on $\operatorname{proj_{f}}$’s output and ensuring effective gradient updates. Meanwhile, the $\operatorname{proj_{v}}$ mapping layer is redefined to map the aggregated similarity information back to the feature space. Notably, we adopt a two-layer perceptron for $\operatorname{proj_{v}}$ instead of a linear layer, as experiments show that single-layer designs offer limited performance gains. With these modifications, both $\operatorname{proj_{f}}$ and $\operatorname{proj_{v}}$ are effectively optimized during training, leading to significant performance improvements.
Figure 4: Cluster pooling configurations. (a) In FEC, $\operatorname{proj_{f}}$ only guides pixel selection and does not update, (b) connecting $\operatorname{proj_{f}}$ and $\operatorname{proj_{v}}$ avoids gradient issues but offers limited performance gains, (c) The proposed cluster pooling adopts a two-layer perceptron for $\operatorname{proj_{v}}$, enabling effective training and improved performance.
@@
This section presents a comprehensive evaluation of the proposed CLUENet, conducts visualization analysis and ablation studies [^1] Further experimental details, visualizations, and failure case analyses are available in the supplementary material..
@@
ShuffleNetv2 (x1.5) (ref6) 2.58 0.31 78.39 90.40 1194.72±23.26 1.1
ShuffleNetv2 (x2.0) (ref6) 6.52 0.64 79.63 90.93 1205.17±7.59 1.3
ConvNeXtv2 (A) (ref43) 3.42 0.55 71.15 84.97 1172.04±34.80 1.6
ConvNeXtv2 (F) (ref43) 4.89 0.79 73.14 86.33 1177.66±21.15 1.8
ConvNeXtv2 (N) (ref43) 15.05 2.46 75.04 87.48 1148.19±28.45 2.8
Attention-based method PVTv2 (b0) (ref11) 3.44 0.54 75.34 88.52 1195.32±31.63 1.7
PVTv2 (b1) (ref11) 13.55 2.06 77.57 89.86 1154.58±39.86 2.9
@@
Datasets. CIFAR-100 and Mini-ImageNet are used for the experiment. CIFAR-100 contains 50K training and 10K validation images of size $32\times 32$. Mini-ImageNet is a subset of ImageNet-1K with 100 classes and 600 images per class. We split Mini-ImageNet 4:1 into training and validation sets, ensuring balanced class distribution.
We follow the standard data augmentation procedure during training. For CIFAR-100, random horizontal flip and normalization are applied. For Mini-ImageNet, random crop, horizontal flip, color jitter, random erasing, and normalization are applied for training.
Training Details. All models are trained using the AdamW optimizer with momentum 0.9 and cosine decay. Weight decay is set to 0.05. The learning rate is auto-tuned using findLR (ref42), with a warmup of 5 epochs. The batch size for CIFAR-100 is 256, and the batch size for Mini-ImageNet is 128. All experiments are conducted on a computer with an NVIDIA RTX 4090 (24GB) GPU.
Evaluation Protocol. Following common practice (ref16; ref19; ref21), all models are evaluated on the validation set after completing 100 training epochs, rather than using the model with the best validation performance, to ensure a fair comparison. Each model is evaluated 7 times; the first 2 runs are discarded to avoid cold-start effects. We report the average of the last 5 runs for classification accuracy (Top-1 and Top-3), throughput (Frames Per Second, FPS), number of parameters (#Param), FLOPs, and inference memory usage, while all metrics except FPS show no significant change.